OpenStudy (anonymous):

An outside circular ring has a circumference of 90 cm. What is the circumference of an inner ring which is 8 cm shorter in radius? Both circles have the same center.

OpenStudy (pfenn1):

The circumference of the out circle is given by what equation?

OpenStudy (pfenn1):

*outer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Let R be the radius of the outer ring. so, 90 = 2piR --> R = 90/(2pi) the circumference of the inner circle is therefore: C = 2pi(R-8) = 2pi[90/(2pi) - 8]

OpenStudy (anonymous):

thank u but that looks in form i dont get

OpenStudy (pfenn1):

If R is the radius of the outer ring, we know the radius of the inner ring is R-8. The circumference of the outer ring is given (90 cm), the formula of the circumference of the outer circle is given above by @dpaInc as 2pi R=90. And @dpalnc shows that if you solve for R you get R=90/2pi.

OpenStudy (pfenn1):

The circumference of the inner circle is given by C=2pi r where r=R-8. So you plug that into the equation and get\[C=2pi (R-8)\] and then plug the above in for R and you get\[C=2pi \left\{ \left( 90 \over 2 pi \right)+8 \right\}\] which is exactly what @dpaInc showed you above. Can you simplify that expression further?

OpenStudy (anonymous):

no but thanxs

OpenStudy (pfenn1):

\[C=2pi \left\{ \left( 90 \over 2 pi \right)-8 \right\}\] Okay I made a sign error above. This is what the equation should look like. It can be simplified to\[C=90-16pi\]